Two masses are placed on the x axis. The first mass 360 kg is located at x = 3.70 cm , the second mass 690 kg is at the origin

Find the gravitational force between the two masses.

3 sig. fig
Answer #1

The gravitational force between the two mass is

F=G\frac{m_1m_2}{r^2}=6.67\times10^{-11}\times\frac{360\times690}{(3.7\times10^{-2}-0)^2}=0.0121N